OpenStudy (anonymous):

Consider the function below. f(t) = 180000-9000t +t^2 Compute and simplify the difference quotient of the given function. f(t+h) - f(t) / all divided (h)

OpenStudy (anonymous):

f(t+h) = 180000-9000(t+h) + (t+h)^2

OpenStudy (anonymous):

So, rest is expanding terms, doing arithmetic, etc.!

OpenStudy (anonymous):

can you show me step by step , im really lost and have tried this problem many times already

OpenStudy (anonymous):

f(t+h) = 180000 - 90000t - 90000h + t^2 + 2th + h^2 f(t) = 180000 - 90000t + t^2 So, f(t+h) - f(t) = 180000 - 90000t - 90000h + t^2 + 2th + h^2 - 180000 + 90000t - t^2 So, f(t+h) - f(t) = 2th + h^2 - 90000h That thing divided by h: 2t + h - 90000

OpenStudy (anonymous):

ummm so whats the answer then?

OpenStudy (anonymous):

I guess that is it.